HETAV PATEL

Grade 11,

a closed pipe and open pipe have their first overtone of same frequency their lenght are in ratio A)1:2B)2:3C)3:4D)4:3

a closed pipe and open pipe have their first overtone of same frequency their lenght are in ratio A)1:2B)2:3C)3:4D)4:3

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
It is given that
First overtone of closed pipe = first overtone of open pipe
3(v/4*l1) = 2(v/2* l2)
Hence
L1 /L2 = ¾
Regards
Arun (askIITians forum expert)